two identical ball dropped from 2000m at differ 2 sec
(1) find the speed of 1st ball will hit ground
(2)assume the height of two points of dropping the two both ball will collide
 

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

The second part is not clear. Please find the answer to the first question below-

Displacement of the ball=s=-2000m
Acceleration of the ball=g=-10m/s^2
Hence from equation of motion,

v^2-u^2=2gs
\implies v^2-0^2=2\times -10\times -2000=40000
\implies v=200m/s